# Visibly Fair Mechanisms

### v1

@@ 5 Achieving Distributional Objectives
When using direct mechanisms, our earlier discussion shows that only serial dictatorship (SD) achieves visible fairness. By eliciting less information about preferences—through carefully designed restricted message spaces—we expand the set of allocations that can be deemed visibly fair for a given problem. This relaxation provides the mechanism designer with additional flexibility, allowing for the implementation of a broader array of allocation rules that still adhere to this notion of fairness.
@@ 5.1 Modular Upper-Bounds
We extend our original model by saying that each officer has a type $t$ from a finite set of types $T$. The distributional goals of the designer are modeled through type-specific modular upper-bounds, where for each type $t\in T$, a collection of upper-bounds specifies limits on the allocation of type-$t$ officers to subsets of states.
@@
Given a finite set of types $T$, for each type $t\in T$, a modular upper-bound system for type $t$, $H^{t}$, is a collection of pairs $h=(S_{h}^{t},k_{h}^{t})$, where:
  1. i)
    $S_{h}^{t}\subseteq S$ is a subset of states,
  2. ii)
    $k_{h}^{t}\in\mathbb{N}$ is an upper-bound on the number of type-$t$ officers that can be assigned to states in $S_{h}^{t}$.
We denote by $H=\left(H^{t}\right)_{t\in T}$ a modular upper-bound system.
@@
An allocation $a\in\mathcal{A}$ respects the modular upper-bounds $H=\left(\right)_{t\in T}$ if, for every type $t\in T$ and every $(S_{h}^{t},k_{h}^{t})\in H^{t}$, the number of type-$t$ officers assigned to states in $S_{h}^{t}$ does not exceed $k_{h}^{t}$, that is,
|$\left|\left\{i\in I:a_{i}\in S_{h}^{t}\text{ and }t_{i}=t\right\}\right|\leq k_{h}^{t}.$
We say that modular upper-bound $(S_{h}^{t},k_{h}^{t})$ is binding at allocation $a\in\mathcal{A}$ if:
|$\left|\left\{i\in I:a_{i}\in S_{h}^{t}\text{ and }t_{i}=t\right\}\right|=k_{h}^{t}$
The literature has proposed several ways to formalise “quota–type” constraints. [^11] An incomplete list includes Echenique and Yenmez (2015), Kamada and Kojima (2015), Goto et al. (2017), Kamada and Kojima (2018), Aziz et al. (2019), Kojima et al. (2020), and Kamada and Kojima (2024). An example of a very permissive notion is the hereditary family of Goto et al. (2017): write a matching as a vector that counts, for every state, how many officers of each type are assigned there; a subset of vectors is hereditary when it is closed under coordinate-wise decreases. Any system of pure ceilings clearly has this property, so every modular upper-bound instance fits inside the hereditary domain. A much tighter specification is the hierarchical (laminar) system analysed by (Kamada and Kojima, 2015; Kamada and Kojima, 2018): here the subsets that carry quotas must form a tree—any two are either disjoint or one contains the other. Laminar caps are useful when the policy maker wants, say, regional ceilings that line up neatly with district ceilings, but they rule out overlapping constraints such as “no more than ten officers in the Northeastern states and no more than eight in the coastal states.” All laminar systems are modular, yet the converse is false.
@@ 5.2 Modular-induced Message Spaces
For each state $s\in S$ and type $t\in T$, define its upper-bound signature as the collection of type-specific upper-bounds that involve $s$ for type $t$:
$H(s,t)=\left\{h:\ (S_{h}^{t},k_{h}^{t})\in H^{t},\ s\in S_{h}^{t}\right\}.$
Definition 13.
For any type $t\in T$, define an equivalence relation tsubscriptsimilar-$\sim_{t}$ on $S$ such that, for all $s,s^{\prime}\in S$,
$s\sim_{t}s^{\prime}\quad\text{if and only if}\quad H(s,t)=H(s^{\prime},t).$
The Modular-induced Message Space associated with the modular upper-bounds $H^{t}$ is the zonal message space $M_{t}$ with zones $Z=\{Z_{1},Z_{2},\dots,Z_{k}\}$, where each zone $Z_{j}$ is an equivalence class under tsubscriptsimilar-$\sim_{t}$, that is,
$Z_{j}=\{s\in S:s\sim_{t}s_{j}\},$
@@
    Zones are disjoint and partition the set of states: $\bigcup_{j}Z_{j}=S$ and $Z_{j}\cap Z_{j^{\prime}}=\varnothing$ for $j\neq j^{\prime}$.
@@
Example 5.1.
  • For type $t_{1}$, the upper-bound system for $t_{1}$ is
[…]
  • For type $t_{2}$, the upper-bound system is
[…]
@@ 5.3 Modular Prioritized Allocation Mechanism
Fix a finite set of agents $I$ and a strict priority order $\pi$ on $I$. Let $H=\bigl{(}H^{t}\bigr{)}_{t\in T}$ be a given modular upper-bound system for a finite set of types $T$. Each type $t\in T$ induces a modular-induced message space $M_{t}$, as in Definition 13, where for each type $t\in T$, $S$ is partitioned into zones according to equivalence classes of states under the same upper-bound signature.
Thus, for an officer $i$ of type $t_{i}$, the mechanism offers the zonal space $M_{t_{i}}$, with associated zones $z_{1},\dots,z_{K}$, requiring her to submit a message $m_{i}\in M_{t_{i}}$. That message ranks all states within each zone but cannot compare states across different zones.
@@
$z_{1}\,\blacktriangleright_{i}\,z_{2}\,\blacktriangleright_{i}\,\cdots\,\blacktriangleright_{i}\,z_{K}$
@@
Given a strict priority ranking $\pi$, a collection of modular upper-bounds $H=\bigl{(}H^{t}\bigr{)}_{t\in T}$, an exogenous zone ranking $\blacktriangleright_{i}$ for each officer $i$, and a profile of messages $m=(m_{i})_{i\in I}$, each $m_{i}$ in the modular-induced message space $M_{t_{i}}$, the Modular Priority Mechanism $\psi$ proceeds as follows:
@@
  • For each $s\in S$, set remaining capacity $C_{s}=q_{s}$.
  • For each zone $z$ and type $t$, set a flag $B_{z}^{t}=\mathrm{False}$.
@@
Sequential Assignment: Process the officers in the order $(i_{1},\dots,i_{n})$ where $\pi(i_{1})<\cdots<\pi(i_{n})$. For each $k=1,\dots,n$:
@@
    • There is a state $s\in z_{\ell}$ for which $C_{s}>0$.
@@
    • Let $s^{*}$ be the most-preferred state of $i_{k}$ within $z_{\ell}$ (according to $m_{i_{k}}$) that still has $C_{s^{*}}>0$.
    • Set $a_{i_{k}}=s^{*}$ and reduce capacity $C_{s^{*}}\leftarrow C_{s^{*}}-1$.
    • For each modular upper-bound $\bigl{(}S_{h}^{t_{i_{k}}},k_{h}^{t_{i_{k}}}\bigr{)}\in H^{t_{i_{k}}}$:
      • If $s^{*}\in S_{h}^{t_{i_{k}}}$, then increase the count of type-$t_{i_{k}}$ officers in $S_{h}^{t_{i_{k}}}$.
      • If this count has now reached $k_{h}^{t_{i_{k}}}$ (i.e. the upper-bound is binding), then for all zones $z^{\prime}\in\{z_{1},\dots,z_{K}\}$ with $z^{\prime}\cap S_{h}^{t_{i_{k}}}\neq\varnothing$, set $B_{z^{\prime}}^{t_{i_{k}}}=\mathrm{True}$.
Outcome: After processing all agents $i_{1},\dots,i_{n}$, the mechanism outputs the allocation $a=(a_{i})_{i\in I}$. We denote
$\psi(m)\;=\;a.$
The Modular Priority Mechanism can be viewed as a partitioned priority mechanism in which each officer type is associated with a zonal message space of states that share the same upper-bound constraints. An officer’s final assignment is determined by two key factors: (1) a counter that tracks remaining capacity for the relevant states, and (2) a flag indicating whether any upper-bound restrictions in that zone have become binding. Because all states in a given zone are governed by the same set of constraints, a single triggered bound applies uniformly across the entire zone. Below, we present the main result and two examples that illustrate how this mechanism operates in practice.
@@
The Modular Priority Mechanism is visibly fair, strategy-proof and respects modular upper-bounds.
@@
Consider four states $S=\{s_{1},s_{2},s_{3},s_{4}\}$, which are partitioned administratively into two regions: $R_{1}=\{s_{1},s_{2}\}\text{ and }R_{2}=\{s_{3},s_{4}\}$, and each $s\in S$ having capacity $q_{s}=2$. There are 8 officers $I=\{i_{1},\dots,i_{8}\}$, and an officer type is the region they are originally from: $t_{i}\in\{R_{1},R_{2}\}$.
The policy objective is that at most 50% of the jobs in a region are taken by local officers. For each $r\in\{R_{1},R_{2}\}$, the modular upper-bound is, therefore:
$H^{r}=\{(R_{r},2)\},$
@@
$H(s,r)=\begin{cases}\{(r,2)\},&\text{if }s\in r,\\[2.84526pt] \varnothing,&\text{if }s\in\hat{r}.\end{cases}$
@@
$z_{1}=\{s_{1},s_{2}\},\quad z_{2}=\{s_{3},s_{4}\}.$
Officers therefore use the same zonal message space, where they submit complete rankings over states within each zone (without comparing states across $z_{1}$ and $z_{2}$).
@@
$\begin{array}[]{rl}z_{1}\;\blacktriangleright_{i}\;z_{2}&\text{if }t_{i}=R_{1}\\ z_{2}\;\blacktriangleright_{i}\;z_{1}&\text{if }t_{i}=R_{2}\end{array}$
@@
An allocation $a$ is feasible if for every $s\in S$, |$|\{i:a_{i}=s\}|\leq q_{s}$, and for each $r\in\{R_{1},R_{2}\}$, |$|\{i:t_{i}=r\text{ and }a_{i}\in R_{r}\}|\leq 2$.
A modular–priority mechanism enforcing these bounds proceeds as in Definition 14, using the zonal message spaces $M_{R_{1}}$ and $M_{R_{2}}$ and exogenous rankings over zones $\left(\blacktriangleright_{i}\right)_{i\in I}$.
@@
with each hospital having capacity $q_{s}=4$. In each $R_{i}$, the first hospital is rural (e.g. $s_{1},s_{4},s_{7}$) and the others are urban ($s_{2},s_{3},s_{5},s_{6},s_{8},s_{9}$). There are 27 doctors, 9 of each type $t\in\{1,2,3\}$, indicating their home region.
@@
$H^{t}=\{(R_{t},6)\}\quad\text{for each }t=1,2,3,\quad\text{and}\quad H^{\mathrm{U}}=\{(U,18)\}$
where $U=\{s_{2},s_{3},s_{5},s_{6},s_{8},s_{9}\}$ is the set of all urban hospitals, and $(U,18)$ is understood to be a universal cap across all doctors.
@@
$H(s,1)=\begin{cases}\{(R_{1},6)\},&\text{if $s\in R_{1}$ is rural (here, $s_{1}$)},\\ \{(R_{1},6),(U,18)\},&\text{if $s\in R_{1}$ is urban (here, $s_{2},s_{3}$)},\\ \{(U,18)\},&\text{if $s$ is urban but not in $R_{1}$ (e.g.\ $s_{5},s_{6},s_{8},s_{9}$)},\\ \varnothing,&\text{if $s$ is rural outside $R_{1}$ (e.g.\ $s_{4},s_{7}$).}\end{cases}$
@@
$Z_{1}=\{s_{1}\},\quad Z_{2}=\{s_{2},s_{3}\},\quad Z_{3}=\{s_{4},s_{7}\},\quad Z_{4}=\{s_{5},s_{6},s_{8},s_{9}\}.$
@@
$Z_{2}\;\blacktriangleright_{i}\;Z_{1}\blacktriangleright_{i}\;Z_{4}\blacktriangleright_{i}\;Z_{3},$
@@
Any allocation must respect both hospital capacities and these upper‐bounds:
|$\bigl{|}\{i:a_{i}=s\}\bigr{|}\leq 4,\quad\bigl{|}\{i:t_{i}=t,\;a_{i}\in R_{t}\}\bigr{|}\leq 6,\quad\bigl{|}\{i:a_{i}\in U\}\bigr{|}\leq 18.$
A modular‐priority mechanism enforces these caps by assigning each $t$ doctor to her top feasible hospital within the partition induced by $(R_{t},6)$ and $(U,18)$.
@@
It is also worth noting that the mechanisms introduced above—and the overarching notion of visible fairness—are ill-suited to implementing within-state affirmative action quotas. Such policies, which often appear in the matching literature on diversity (e.g. majority quotas or type-specific ceilings in each state), could be encoded as modular upper-bounds with one state in each upper-bound. However, a scenario in which, for example, each state can admit a maximum number of agents of a certain type, the resulting zonal message space would have one state per zone. This, of course, would completely eliminate role of preferences in the allocation. If we attempted to compensate this by using zonal message spaces with ranking over zones, we would easily conclude that there is no visibly fair mechanism with these characteristics that would respect upper bounds.

### v2

@@ 5 Achieving Distributional Objectives
When using direct mechanisms, our earlier discussion shows that only serial dictatorship (SD) achieves visible fairness. By eliciting less information about preferences—through carefully designed restricted message spaces—we expand the set of allocations that can be deemed visibly fair for a given problem. This relaxation provides the policy maker with additional flexibility, allowing for the implementation of a broader array of allocation rules that still adhere to this notion of fairness.
@@ 5.1 Modular Upper-Bounds
We extend our original model by saying that each officer $i$ has a type $t$ from a finite set of types $T$, where $t_{i}$ denotes the type of officer $i$. The distributional goals of the designer are modeled through type-specific modular upper-bounds, where for a set of types, a collection of upper-bounds specifies limits on the allocation of officers with those types to subsets of states.
@@
A modular upper‑bound system is a finite collection
$H\;=\;\bigl{\{}\,(\Xi_{h},\;S_{h},\;k_{h})\bigr{\}},$
where for each element $h\in H$: [^15] In our notation, $H$ contains a set of upper-bounds, each of which represented by the letter $h$. In this context, $\Xi_{h}$ for example, is the first component of $h$.
  • $\varnothing\neq\Xi_{h}\subseteq T$ is the set of types covered by the quota,
  • $S_{h}\subseteq S$ is a subset of states, and
  • $k_{h}\in\mathbb{N}$ is the ceiling.
For every type $t\in T$ we write
$H^{t}\;:=\;\bigl{\{}\,(\Xi_{h},S_{h},k_{h})\in H:t\in\Xi_{h}\bigr{\}}.$
For every state $s\in S$ and type $t\in T$, we write its upper-bound signature as
$H^{s,t}=\{\,(\Xi_{h},S_{h},k_{h})\in H:s\in S_{h},\;t\in\Xi_{h}\}.$
@@
An allocation $a\in\mathcal{A}$ respects the modular upper-bounds $H$ if, for every $(\Xi_{h},S_{h},k_{h})\in H$:
$\left|\{\,i\in I:\;a_{i}\in S_{h},\;t_{i}\in\Xi_{h}\}\right|\;\leq\;k_{h}.$
We say that the modular upper-bound $(\Xi_{h},S_{h},k_{h})$ is binding at allocation $a\in\mathcal{A}$ if the constraint is satisfied with equality.
Since in our model officers cannot be left unmatched, we need to guarantee that these upper-bounds are compatible with that restriction while using visibly fair mechanisms. Formally, hereafter we will restrict our attention to modular upper-bound systems that satisfy the following property of sequential solvency.
Definition 12.
For any allocation $a\in\mathcal{A}$, officer $i\in I$, and modular upper bound $h\in H$, define $n_{s}\;:=\;\bigl{|}\{j\neq i:a_{j}=s\}\bigr{|}$ and $n_{h}\;:=\;\bigl{|}\{j\neq i:a_{j}\in S_{h},\;t_{j}\in\Xi_{h}\bigr{|}$. The modular upper-bounds system $H$ satisfies sequential solvency if and only if for all $i\in I$ and for all $a\in\mathcal{A}$ that respect $H$,
$\exists s\in S\;\text{s.t. }\;\begin{cases}n_{s}<q_{s},\text{ and }\\[4.0pt] h\in H^{s,t_{i}}\ \Longrightarrow\ n_{h}<k_{h}\quad\forall h\in H.\end{cases}$
The intuition for the definition above is simple. Regardless of which capacities or collection of upper-bounds bind, there will always be a compatible state for every remaining officer. That is, while matching officers one at a time, modular upper-bounds can restrict where officers are matched, but not whether they are matched. Since it relies on the particular number of agents of each type in $I$, it allows for interesting and practical constraints, as we will show in examples that will follow. [^16] Notice that it is crucial that the definition depends on the profile of types of officers. Otherwise, the upper-bounds would have to be satisfied when all agents have the same type, making only bounds that never bind compatible with not leaving officers unmatched.
Example 5.1.
  • For type $t_{1}$, the upper-bound system is
    $H^{t_{1}}=\Bigl{\{}\bigl{(}\{t_{1}\},\{s_{1},s_{2},s_{3}\},\,2\bigr{)}\Bigr{\}},$
  • For type $t_{2}$, the upper-bound system is
    $H^{t_{2}}=\Bigl{\{}\bigl{(}\{t_{2}\},\{s_{3},s_{4},s_{5}\},\,1\bigr{)}\Bigr{\}},$
The literature has proposed several ways to formalize “quota–type” constraints. [^17] An incomplete list includes Echenique and Yenmez (2015), Kamada and Kojima (2015), Goto et al. (2017), Kamada and Kojima (2018), Aziz et al. (2019), Kojima et al. (2020), and Kamada and Kojima (2024). An example of a very permissive notion is the hereditary family of Goto et al. (2017): write a matching as a vector that counts, for every state, how many officers of each type are assigned there; a subset of vectors is hereditary when it is closed under coordinate-wise decrements. Any system of pure ceilings clearly has this property, so every modular upper-bound instance fits inside the hereditary domain. A tighter specification is the hierarchical (laminar) system analysed by (Kamada and Kojima, 2015; Kamada and Kojima, 2018): here the subsets that carry quotas must form a tree—any two are either disjoint or one contains the other. Laminar caps are useful when the policy maker wants, say, regional ceilings that line up neatly with district ceilings, but they rule out overlapping constraints such as “no more than ten officers in the Northeastern states and no more than eight in the coastal states.” All laminar systems are modular, yet the converse is false.
@@ 5.2 Modular-induced Message Spaces
For any type $t\in T$, define an equivalence relation $\sim_{t}$ on $S$ such that, for all $s,s^{\prime}\in S$,
$s\sim_{t}s^{\prime}\quad\text{if and only if}\quad H^{s,t}=H^{s^{\prime},t}.$
The Modular-induced Message Space associated with the modular upper-bounds $H^{t}$ is the zonal message space $M_{t}$ with zones $Z=\{z^{t}_{1},z^{t}_{2},\dots,z^{t}_{k}\}$, where each zone $z^{t}_{j}$ is an equivalence class under $\sim_{t}$, that is,
$z^{t}_{j}=\{s\in S:s\sim_{t}s_{j}\},$
@@
    Zones are disjoint and partition the set of states: $\bigcup_{j}z^{t}_{j}=S$ and $z^{t}_{j}\cap z^{t}_{j^{\prime}}=\varnothing$ for $j\neq j^{\prime}$.
@@ 5.3 Modular Prioritized Allocation Mechanism
Fix a finite set of agents $I$ and a strict priority order $\pi$ on $I$. Let $H$ be a given modular upper-bound system for a finite set of types $T$. Each type $t\in T$ induces a modular-induced message space $M_{t}$, as in Definition 13, where for each type $t\in T$, $S$ is partitioned into zones $Z^{t}=\{z^{t}_{1},z^{t}_{2},\ldots\}$ according to equivalence classes of states under the same upper-bound signature.
Thus, for an officer $i$ of type $t_{i}$, the mechanism offers the zonal space $M_{t_{i}}$, with associated zones $z^{t_{i}}_{1},z^{t_{i}}_{2}\ldots$, requiring her to submit a message $m_{i}\in M_{t_{i}}$. That message ranks all states within each zone but cannot compare states across different zones.
@@
$z^{t_{i}}_{1}\,\blacktriangleright_{i}\,z^{t_{i}}_{2}\,\blacktriangleright_{i}\,\cdots\,\blacktriangleright_{i}\,\ldots$
@@
Given a strict priority ranking $\pi$, a modular upper-bound system $H$, an exogenous zone ranking $\blacktriangleright_{i}$ for each officer $i$, and a profile of messages $m=(m_{i})_{i\in I}$, each $m_{i}$ in the modular-induced message space $M_{t_{i}}$, the Modular Priority Mechanism $\psi$ proceeds as follows:
@@
  • For each $s\in S$, set remaining capacity $q_{s}^{rem}=q_{s}$.
  • For each type $t$ and zone $z^{t}_{\ell}$, set a flag $B_{\ell}^{t}=\mathrm{False}$.
@@
Quotas update procedure:
We next describe the procedure that updates the zone flags below:
  • For each $(\Xi_{h},S_{h},k_{h})\in H$:
    • Let $N_{h}=\left|\left\{i\in I:a_{i}\in S_{h}\text{ and }t_{i}\in\Xi_{h}\right\}\right|$.
    • If $N_{h}=k_{h}$, then for every $t\in\Xi_{h}$ and $\ell$ such that $S_{h}\cap z^{t}_{\ell}\neq\varnothing$, set $B_{\ell}^{t}=\mathrm{True}$.
Sequential Assignment: Process the officers in the order $(i_{1},\dots,i_{n})$. For each $k=1,\dots,n$:
@@
    • There is a state $s\in z_{\ell}$ for which $q_{s}^{rem}>0$.
@@
    • Let $s^{*}$ be the most-preferred state of $i_{k}$ within $z_{\ell}$ (according to $m_{i_{k}}$) that still has $q_{s^{*}}^{rem}>0$.
    • Set $a_{i_{k}}=s^{*}$ and reduce capacity $q_{s^{*}}^{rem}\leftarrow q_{s^{*}}^{rem}-1$.
    • Apply the quotas update procedure.
Outcome: After processing all agents $i_{1},\dots,i_{n}$, the mechanism outputs the allocation $a=(a_{i})_{i\in I}$.
The Modular Priority Mechanism is, therefore, a partitioned priority mechanism in which each officer type is associated with a zonal message space of states that share the same upper-bound constraints. An officer’s final assignment is determined by two key factors: (1) a counter that tracks remaining capacity for the relevant states, and (2) a flag indicating whether any upper-bound restrictions in that zone have become binding. Because all states in a given zone are governed by the same set of constraints, a single triggered bound applies uniformly across the entire zone. Below, we present the main result for this mechanism and two examples that illustrate how it operates in practice. [^18] In Appendix B we show that the modular upper-bounds system presented in Example 5.2 and Example 5.3 satisfy sequential solvency.
@@
The Modular Priority Mechanism is visibly fair, strategy-proof, and respects modular upper-bounds.
@@
Consider four states $S=\{s_{1},s_{2},s_{3},s_{4}\}$, which are partitioned administratively into two regions: $R_{1}=\{s_{1},s_{2}\}\text{ and }R_{2}=\{s_{3},s_{4}\}$, and each $s\in S$ having capacity $q_{s}=2$. There are 8 officers $I=\{i_{1},\dots,i_{8}\}$, and an officer type corresponds to the region they are originally from: $t_{i}\in\{1,2\}$.
The policy objective is that at most 50% of the jobs in a region are taken by local officers. For each $r\in\{1,2\}$, the modular upper-bound is, therefore:
$H^{r}=\{(\{r\},R_{r},2)\},$
@@
[…]
@@
$z^{r}_{1}=\{s_{1},s_{2}\},\quad z^{r}_{2}=\{s_{3},s_{4}\}.$
Officers therefore use the same zonal message space, where they submit complete rankings over states within each zone (without comparing states across $z^{r}_{1}$ and $z^{r}_{2}$).
@@
$\begin{array}[]{rl}z^{r}_{1}\;\blacktriangleright_{i}\;z^{r}_{2}&\text{if }t_{i}=1\\ z^{r}_{2}\;\blacktriangleright_{i}\;z^{r}_{1}&\text{if }t_{i}=2\end{array}$
@@
A modular–priority mechanism enforcing these bounds proceeds as in Definition 14, using the zonal message spaces $M_{1}$ and $M_{2}$ and exogenous rankings over zones $\left(\blacktriangleright_{i}\right)_{i\in I}$.
@@
with each hospital having capacity $q_{s}=4$. In each region, the first hospital is rural ($s_{1},s_{4},s_{7}$) and the others are urban ($s_{2},s_{3},s_{5},s_{6},s_{8},s_{9}$). There are 27 doctors, 9 of each type $t\in\{1,2,3\}$, indicating their home region.
@@
$H^{t}=\{(\{t\},R_{t},6)\}\quad\text{for each }t=1,2,3,\quad\text{and}\quad H^{\mathrm{U}}=\{(\{1,2,3\},U,19)\}$
where $U=\{s_{2},s_{3},s_{5},s_{6},s_{8},s_{9}\}$ is the set of all urban hospitals, and $(U,19)$ is a universal cap across all doctors.
@@
[…]
@@
$z^{1}_{1}=\{s_{1}\},\quad z^{1}_{2}=\{s_{2},s_{3}\},\quad z^{1}_{3}=\{s_{4},s_{7}\},\quad z^{1}_{4}=\{s_{5},s_{6},s_{8},s_{9}\}.$
@@
$z^{1}_{2}\;\blacktriangleright_{i}\;z^{1}_{1}\blacktriangleright_{i}\;z^{1}_{4}\blacktriangleright_{i}\;z^{1}_{3},$
@@
A modular‐priority mechanism enforces these caps by assigning each type‐$t$ doctor to her top feasible and non-binding hospital within the partition induced by $(R_{t},6)$ and $(U,19)$.
@@
It is also worth noting that the mechanisms introduced above—and the overarching notion of visible fairness—are ill-suited to implementing within-state affirmative action quotas. Such policies, which often appear in the matching literature on diversity (e.g. majority quotas or type-specific ceilings in each state), could be encoded as modular upper-bounds with one state in each upper-bound. However, a scenario in which, for example, each state can admit a maximum number of agents of a certain type, the resulting zonal message space would have one state per zone. This, of course, would eliminate a role for preferences in the allocation. If we attempted to compensate for this by using zonal message spaces with ranking over zones, we would easily conclude that no visibly fair mechanism with these characteristics would respect upper bounds.